OpenStudy (nicoleced):

(2a-3b+5c) - (7a+9b+6c) I think it's -5a^2 -6b^2 -1c^2 wrong, or right? please correct me and explain, thanks!

OpenStudy (anonymous):

right

OpenStudy (anonymous):

is this multiplying or subtrcting..???

OpenStudy (nicoleced):

subtracting. sorry

OpenStudy (anonymous):

then its wrong..!!!

OpenStudy (nicoleced):

okay, so how do I do it?

OpenStudy (anonymous):

to subtract simply subtract the coefficients of like terms..!!!e'g' (2-7)a+....

OpenStudy (nicoleced):

yeah, that's what I did... no?

OpenStudy (anonymous):

you did thiis but you multiply the variables too..!!!you donot have to do that..

OpenStudy (nicoleced):

2-7 -5a^2

OpenStudy (anonymous):

subtract the coefficeints by taking variable as common...not to multiply...

OpenStudy (nicoleced):

so the answer would just be 5a-6b-1c?

OpenStudy (anonymous):

see it below.. (2-7)a+(-3-9)b+(5-6)c now solve it for answer..

OpenStudy (nicoleced):

-5a-6b-1c ... right?

OpenStudy (nicoleced):

@zaibali.qasmi

OpenStudy (anonymous):

do it again..still wrong ...you did some mistake..!!

OpenStudy (nicoleced):

what mistake? I don't understand what I'm doing wrong.

OpenStudy (nicoleced):

@zaibali.qasmi

OpenStudy (anonymous):

-3-9= what..???

OpenStudy (nicoleced):

-6

OpenStudy (anonymous):

its your mistake... -3-9=-12 not -6

OpenStudy (nicoleced):

why? I get -6

OpenStudy (anonymous):

no..-3-9=-12 coz similar signs always give rise to additio with same sign..

OpenStudy (nicoleced):

that crap still gets me confused. ugh

OpenStudy (anonymous):

but why..do you noot know the addition and subtraction..??

OpenStudy (nicoleced):

of course I know how to add and subtract, its just the negatives and positives that I still don't get. and I have my final in a few hours. I'm fluttered!